OpenStudy (jessie13):

If the dimensions of the following cylinder are doubled, what will be the surface area of the new cylinder?

OpenStudy (anonymous):

welcome to freakin open study

OpenStudy (rhystic):

they are quadrupled because the area of a cyclinder is given by 2*pi*r^2 + 2*pi*r*h

OpenStudy (jessie13):

I still don't get it!

OpenStudy (rhystic):

so its because you have to square something to get the area. squaring means multiplying the value by itself. If you double the value before squaring then it yields a number 4 times that of the inital value squared. E.g.: 2 ; 2^2 = 4 double 2 and you get 4 so 4^2 = 16 ; or lets go with 6 : 6^2 = 36 double six you get 12 : 12^2 = 144 36*4 is 144 so regardless of the number when you double it and have to square that number youve effectively quadrupled the area.

OpenStudy (jessie13):

so 144 is the answer

OpenStudy (rhystic):

what? no. you didn't give any dimensions of the cylinder all I am saying is doubling the dimensions will quadruple the area of the cylinder. If you have numbers and need to solve something you need to post that...